The fireworks shoot off to the east from the launchpad at an angle of 65 with the ground. Elora is standing 290 feet to the west of the launchpad. Elora sees the fireworks at an angle of elevation of 40. What is the height of the fireworks when they explode
Answer by mananth(16946)

.
tan 65 = h/x
h = xtan65
tan 40 = h/(x+290)
h= (x+290) *tan40
Therefore
x*tan65= (x+290) *tan40
x*tan65= x*tan40 + 290 *tan 40
x*tan65-x*tan40 = 290 *tan 40
x(tan 65 - tan 40 ) = 243.34
1.30x = 243.4
x= 243.4/1.3
x=187.23
tan 65 = h/187.23
h= 401.5 ft
height of explosion